# MetAdv: A Unified and Interactive Adversarial Testing Platform for Autonomous Driving

### v2

@@ Abstract.
Autonomous driving, Adversarial attacks, Simulation sandbox
@@
journalyear: 2025ccs: Security and privacyccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning
@@ 1. Introduction
Autonomous driving has made significant progress in recent years (Zimmer et al., 2019; Lang et al., 2019; Erçelik et al., 2022; Zimmer et al., 2022; Li et al., 2024; Hu et al., 2023b; Chen et al., 2024; Jiang et al., 2023), yet remain highly susceptible to adversarial attacks (Szegedy et al., 2013; Goodfellow et al., 2014; Liang et al., 2020; Wei et al., 2018; Liang et al., 2022b, a; Kong et al., 2024), ranging from digital perturbations in sensor input (Dong et al., 2023; Hallyburton et al., 2022; Liu et al., 2020b; Zhang et al., 2021) to physical-world camouflage (Eykholt et al., 2018; Liu et al., 2019; Wang et al., 2021; Liu et al., 2021, 2020a, 2023a, 2022, 2023c; Guo et al., 2023; Liu et al., 2023b; Zhang et al., 2024b; Wang et al., 2024, 2025; Zhang et al., 2024a), which can lead to critical failures and safety hazards. As these systems move closer to widespread deployment, evaluating and ensuring their adversarial robustness has become an urgent and unresolved challenge. However, current adversarial testing approaches are hampered by the lack of a unified, interactive, and dynamic evaluation platform. Existing studies (Xu et al., 2022; Li et al., 2022; Tang et al., 2021) often focus narrowly on isolated and static environments, limiting their capacity to uncover the full spectrum of weaknesses that may emerge in real-world conditions.
@@ 2.1. Adversarial Testing Sandbox
We first propose a unified closed-loop evaluation framework that enables the real-time adversarial testing and bridges the gap between simulation and real-world AD systems; we then propose the dynamic adversarial testing evolution that can generate adversarial test cases targeted to the AD.
@@ 2.2. AD Asset Repository
AD Models and Tasks. MetAdv supports a diverse range of AD algorithms covering perception, prediction, and planning stages, and includes 7 classical AD tasks (e.g., obstacle recognition, scene understanding). Specifically, MetAdv consists of 10 deep learning based perception models (e.g., YOLOs (Redmon et al., 2016), SMOKE (Liu et al., 2020c)), 5 reinforcement learning based decision AD models (e.g., DDPG (Lillicrap et al., 2015), PPO (Schulman et al., 2017)). Additionally, we incorporate unified frameworks that cover the full pipeline, including 4 end-to-end models (e.g., UniAD (Hu et al., 2023a), VAD (Jiang et al., 2023)) and 3 LLM-based AD models (e.g., LMDrive (Shao et al., 2024), Dolphins (Ma et al., 2023)). We also introduce 6 high-quality 3D vehicle models represented in .obj files, such as Tesla Model 3.

### v3

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754477isbn: 979-8-4007-2035-2/2025/10
@@ 1. Introduction
Autonomous driving has made significant progress in recent years (Zimmer et al., 2019; Lang et al., 2019; Erçelik et al., 2022; Zimmer et al., 2022; Li et al., 2024; Hu et al., 2023b; Chen et al., 2024; Jiang et al., 2023), yet remain highly susceptible to adversarial attacks (Szegedy et al., 2013; Goodfellow et al., 2014), ranging from digital perturbations in sensor input (Dong et al., 2023; Hallyburton et al., 2022) to physical-world camouflage (Eykholt et al., 2018; Liu et al., 2019; Wang et al., 2021), which can lead to critical failures and safety hazards. As these systems move closer to widespread deployment, evaluating and ensuring their adversarial robustness has become an urgent and unresolved challenge. However, current adversarial testing approaches are hampered by the lack of a unified, interactive, and dynamic evaluation platform. Existing studies (Xu et al., 2022; Li et al., 2022) often focus narrowly on isolated and static environments, limiting their capacity to uncover the full spectrum of weaknesses that may emerge in real-world conditions.
@@ 2.1. Adversarial Testing Sandbox
We first propose a unified closed-loop evaluation framework that enables the real-time adversarial testing and bridges the gap between simulation and reality; we then propose the dynamic adversarial testing evolution that generate adversarial cases targeted to AD.
@@ 2.2. AD Asset Repository
AD Models and Tasks. MetAdv supports a diverse range of AD algorithms covering perception, prediction, and planning stages, and includes 7 classical AD tasks (e.g., obstacle recognition, scene understanding). Specifically, MetAdv consists of 10 deep learning based perception models (e.g., YOLOs (Redmon et al., 2016), SMOKE (Liu et al., 2020)), 5 reinforcement learning based decision AD models (e.g., DDPG (Lillicrap et al., 2015), PPO (Schulman et al., 2017)). Additionally, we incorporate unified frameworks that cover the full pipeline, including 4 end-to-end models (e.g., UniAD (Hu et al., 2023a), VAD (Jiang et al., 2023)) and 3 LLM-based AD models (e.g., LMDrive (Shao et al., 2024), Dolphins (Ma et al., 2023)). We also introduce 6 high-quality 3D vehicle models represented in .obj files, such as Tesla Model 3.